Saturday 12 6
This presentation is the property of its rightful owner.
Sponsored Links
1 / 9

Saturday 12.6 PowerPoint PPT Presentation


  • 85 Views
  • Uploaded on
  • Presentation posted in: General

Saturday 12.6. Initial plan for the Friday: Morning: Collimator setup at 3.5 TeV . Evening: Transverse feedback system commissioning. Night (if time permits): Squeeze setup with separated beams. 09: 0 0 : Start injection setup 10:00: Injected single bunches with 9.2 10 10 in both beams.

Download Presentation

Saturday 12.6

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Saturday 12 6

Saturday 12.6

  • Initial plan for the Friday:

    • Morning: Collimator setup at 3.5 TeV.

    • Evening: Transverse feedback system commissioning.

    • Night (if time permits): Squeeze setup with separated beams.

  • 09:00 : Start injection setup

    • 10:00: Injected single bunches with 9.2 1010 in both beams.

    • Emittances of 3-4 mm in both planes for both beams.

    • Increased Q’ to 10 units in both planes for both beams following recommendation by E.Metral.

    • Emittances at end of ramp: 6-7 mm in both planes for both beams.

    • 11:30: trimmed all octupoles to 100A.

8:30 meeting


Saturday 12 61

Saturday 12.6

  • 11:45: Start collimation setup.

    • 13:15: Beams dumped due to losses while aligning the TCTH.4R2.B1. Step of 20 microns!

    • Completed 4 collimators per beam  ca. 15 min per collimator.

  • 14:30: Circuit RQ5.L8 tripped:

    • Water problem on the water-cooled cable of B2. Requires access.

    • 16:10: End of access in point 8. CV has clean-up and repositioned the filter on RQ5.L8B2.

    • Flow was noted to be 20l/minute. Filters were cleaned but not very dirty. However, the filter was deformed. After manipulation of the filter, the water flow went up to 35l/minute.

    • 17:20: trip of RCBXV2.L1 during pre-cycle.

    • 18:05: RCBXV2.L1 tripped again with the same fault.

    • beam 2 not extracted because of BIC crates down in SPS.

8:30 meeting


Saturday 12 62

Saturday 12.6

  • 18:15:Pre-Cycle finished:

    • 18:40: Beam 2 not extracted because of BIC crates down in SPS.

    • 19:15: Beam2 is back.

  • 20:30: Starting damper studies at 450 GeV:

    • Both beams with 1 bunches a 2e10.

    • Emittances: B2H 2.4 mm; B2V 3.1 mm.

    • All coarse delays adjusted for all dampers.

    • All phase shifters adjusted, for damping in all planes.

    • Two pick-up mode with vector sum: coarse settings for H.B1, V.B1 and H.B2 done; For V.B2: something has changed with respect to pervious setting-upsessions; Optimum phase setting using Q9 only has changed from 100 degrees to 70 degrees; not understood even after dedicated phase measurements.

8:30 meeting


Saturday 12 63

Saturday 12.6

  • Damper Studies Summary continued:

    • Damper caused 80 % of beam lost during Beam 1 injection. It was traced backto a wrong setting for high intensity.

    • Ramping with all damper modules on and tune feedback is not obvious. Duringthe ramp preparation the damper experts were reducing the damper gain as much as possible and still the tune signal is not very good. The reduction of the gain then caused again some emittance growth, otherwise the emittances stayed small. The tune feedback trims are very jittery.

    • Observed case were the damper could create losses in IR2 (missing the collimators in IR3 and IR7)  to be understood; wrong hierarchy might be related to TDI settings (as discussed 2 days ago: TDI position dependence on primary position).

  • 06:30: End Damper studies and preparing fill for Collimator setup:

    • High intensity fill with 9 1010 single bunch per beam.

8:30 meeting


Tune spectrum with damper on

Tune Spectrum with Damper ON

LMC report


Beam losses with damper on

Beam Losses with Damper ON

LMC report


Trajectory during beam loss

Trajectory during Beam Loss

LMC report


Sunday 13 6

Sunday 13.6

  • Initial plan for the Friday:

    • Morning: Collimator setup at 3.5 TeV.

    • Evening: Transverse feedback system commissioning.

    • Night (if time permits): Squeeze setup with separated beams.

  • Open Issues:

    • RCS.A45B1 QPS not OK; working on board B now.

    • RCBXV2.L1 tripped twice within one hour

8:30 meeting


Program for the next 2 weeks

Program for the next 2 weeks

LMC report


  • Login